. Find the nth term of the following sequences:

a.) 5, 10, 15, 20, …

b.) 22, 18, 14, 10, …


1
Expert's answer
2021-03-16T08:05:09-0400

a) We have an arithmetic progression


"a_1=5, d=10-5=15-10=20-15=5"

"a_n=a_1 +d(n-1)"

"a_n=5+5(n-1), n=1, 2, ..."

"a_n=5n, n=1, 2, ..."

b) We have an arithmetic progression


"a_1=22, d=18-22=14-18=10-14=-4"

"a_n=a_1 +d(n-1)"

"a_n=22+(-4)(n-1), n=1, 2, ..."

"a_n=26-4n, n=1, 2, ..."